Land and a warehouse were acquired for $980,000. What amounts should be recorded in the accounting...

  1. Land and a warehouse were acquired for $980,000. What amounts should be recorded in the accounting records for land and for the warehouse if an appraisal showed the estimated values to be $450,000 for the land and $750,000 for the warehouse?
    1. $450,00 for land; $530,000 for warehouse
    2. $367,500 for land; $612,500 for warehouse
    3. $450,000 for land; $750,000 for warehouse
    4. $230,000 for land; $750,000 for warehouse
  2. On March 2, 2005, Kent industries purchased a fleet of automobiles at a cost of $450,000. The cars are to be depreciated by the straight-line method over five years with no salvage value. Kent uses the half-year convention to compute depreciation for fractional periods. The book value of the fleet of automobiles at December 31,2006 will be:
    1. $135,000
    2. $300,000
    3. $405,000
    4. $315,000
Answer #1

1.

Total value of Land and warehouse = 980000
We will apportionate the value in the ratio of fair value
Land 450000 x980000 367500
1200000
Warehouse 750000 x 980000 612500
1200000
980000

B is correct

2.Book value at year end 2005

Depreciation = 450000 / 5 x 6 /12 = 45000

Book Value = 450000 -45000 = 405000

Book value at year end 2006

Depreciation = 450000 /5 = 90000

Book Value = 405000 -90000 = 315000

D is correct